OpenStudy (anonymous):

solve the fallowing system of equation 3x-5y=4 7x+11y=-2 please!

OpenStudy (anonymous):

7(3x-5y=4) 3(7x+11=-2) 21x-35y=28 21x+33y=-6 21s cancel out leaving -35y=28 minus 33y=-6. Then solve for y. When you get an integer for y plug it in for y in one of the equations then solve for x